Status Report

ISS On-Orbit Status 26 Aug 2002

By SpaceRef Editor
August 26, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below. Start of Week 12 of Expedition Five.

Valery Korzun and Sergey Treschev’s EVA-8 from the DC-1 module was completed successfully, throughout smoothly and in less time than expected. Hatch opening, at 1:27am (all times Eastern daylight), was 25 minutes late because of a temporary air leak between DC-1 and SM Transfer Compartment (PkhO). After repressurization and hatch closure repeat, the spacewalk proceeded nominally, concluding with hatch close at 6:48am, after a total duration of 5 h 21 min. [All five scheduled tasks were completed OK: (1) installation of a platform frame (PF) for temporary ORU (on-orbit replaceable unit) stowage outside the FGB, (2) installation of "fairleads" along the FGB (fixtures for routing and attaching tethers used as translation guides/aids in future EVAs), (3) retrieval of panel #1 of Japanese payloads MPAC (debris particle collector) and SEED (materials exposure experiment), (4) changeout of the Russian Kromka experiment "witness" (collector) plate, designed to collect thruster plume residue to check out the efficacy of plume deflectors installed last year on the SM, and (5) installation of two more ham radio antennas, WA3 and WA4, at the SM aft end, with routing and installation of their cabling.]

Pre-EVA operations after wakeup last night (5:00pm ) proceeded essentially on schedule. [They included: Preparation of DC-1 and PkhO, rerouting C&W (caution and warning) alarms from the SM C&W panel (SPP) to the EVA support panels (POV) in DC-1 and PkhO, deactivation and removal of SM ventilation and air ducts (to free hatches for closure), checkout of the Orlan suits with their comm and medical equipment, EVA gear donning, final tests and airlock depress.]

After Korzun and Treschev’s return shortly before 7:00 am this morning, FE-1 Peggy Whitson assisted in returning the ISS to its initial configuration, i.e., DC-1 air duct installation, replacing the protective cover on the EV hatch, and reactivating ISS systems that had been shut down for the duration of the spacewalk. The CDR dismantled the Glisser-M video system for storage.

Before and after their excursion, Valery and Sergey both completed sessions with the mandatory Russian biomed tests MO-8 (body mass measurement) and MO-9 (urinalysis), which Whitson had set up beforehand. She subsequently performed closeout operations and stowed the experiment hardware.

Station interior was repressurized with N2 (nitrogen) from the Airlock HPGTs (high-pressure gas tanks) to a target pressure of 735 mmHg. The atmosphere was also enriched with fresh O2 (oxygen) from the storage system (SRPK) in Progress 8P.

The ground powered down the SSRMS (space station remote manipulator system), whose video cameras had covered the EVA. The arm and the MBS (mobile base system) were put on Keep-Alive power on both electronics strings, and the Lab RWS (robotic workstation) was turned off. Later, Peggy Whitson also disconnected the RWS’ DCP (display & control panel) power supply cable from the UOP (utility operations panel).

At about 10:32 am, the station’s flight attitude returned to sun-pointing XPOP mode (x-axis perpendicular to orbit plane), torqued by the CMGs at a nearly imperceptible 0.05 deg per sec for 40 minutes.

Sleep time began for the crew today at 12:20 pm, to last until 10:30 pm tonight. Activity schedule for tomorrow features post-EVA tagup with MCC-M, closeout ops for MPAC & SEED with HDTV (high-definition television) coverage, PuFF ops for Korzun and Treschev and a partial crew day off.

Today’s targets for the CEO (crew earth observations) program were Kingman Reef (detailed views are needed for mapping project. Just left of track), Palmyra Atoll (detailed nadir views needed for more accurate mapping of these remote islets), Amman, Jordan (desert city on east side of the Dead Sea, built of country rock, can be difficult to detect. Crew was to try to get the whole built-up area in one frame. Nadir view. ESC [electronic still camera] requested), Guangzhou, China (nadir view, with ESC), Xianggang (Hong Kong), China (nadir view, with ESC), Hurricane Faust (textbook hurricane may have been visible in the distance off to the crew’s right).

CEO images can be viewed at the website

U.S. and Russian Segment Status (as of 10:37 am EDT):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
Elektron O2 generator is powered On (32-amp mode), on backup pump. Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5, i.e., 10-min. cycle time (vacuum pump failed). U.S. CDRA CO2 scrubber is offline. BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Purify mode.

SM Working Compartment: Pressure (mmHg) — 743, temperature (deg C) — 25.7, ppO2 (mmHg) — 154.0, ppCO2 (mmHg) — 2.3.
SM Transfer Compartment: Pressure (mmHg) — 736, temperature (deg C) — 20.2.
FGB Cabin: Pressure (mmHg) — 744, temperature (deg C) — 21.7.
Node: Pressure (mmHg) — 740.71, temperature (deg C) — 22.3 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
U.S. Lab: Pressure (mmHg) — 742.49, temperature (deg C) — 24.8, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
Joint Airlock (Equip. Lock): Pressure (mmHg) — 742.49, temperature (deg C) — 24.8; shell heater temp (deg C) — 25.8, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
PMA-1: Shell heater temp (deg C) — 22.1
PMA-2: Shell heater temp (deg C) — 17.0

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]). MSA (mass spectrometer assembly) and VGA (verification gas assembly) were replaced, but some more work needs to be done).

Electrical Power Systems (EPS):
Both P6 channels fully operational. Beta Gimbal Assembly (BGA) 2B and BGA 4B in Autotrack mode (solar-tracking).
SM batteries: Battery #1 is off-line; all other batteries (7) are in "Full Charge" mode.
FGB batteries: Battery #5 is off-line; all other batteries (5) are in "Partial Charge" mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:
Air conditioner SKV-1 is On. SKV-2 is Off.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1 is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM is Off.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.

Attitude Source:
3 CMGs on-line.
State vector — US GPS (SIGI string 1)
Attitude — Russian segment
Angular rates — US RGA1 (rate gyro assembly 1)

Communications & Tracking Systems:
All Russian communications & tracking systems are nominal.
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is operating nominally.

SSRMS/Canadarm2 at MBS PDGF 1 (mobile base system/power & data grapple fixture 1), in EVA-8 viewing position, with Keep Alive power on both strings (based on MBS).
MBS: Keep Alive power on both strings.
RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

ISS Orbit (as of this morning, 7:31am EDT [= epoch]):
Mean altitude — 391.7 km
Apogee — 402.8 km
Perigee — 380.6 km
Period — 92.4 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0016396
Orbits per 24-hr. day — 15.58
Solar Beta Angle — -41.0 deg (magnitude decreasing)
Altitude decrease — 250 m (mean) in last 24 hours
Revolutions since FGB/Zarya launch (Nov. ’98) — 21499
Current Flight Attitude — XPOP (x-axis perpendicular to orbit plane = "sun-fixed" [yaw: ~180 deg, pitch: -5.7 deg., roll: 0 deg]).

For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.